Find the accumulated amount A if the principal P is invested at the interest rate of r/year for t yr. (Round your answer to the nearest cent.)
P = $1000, r = 4%, t = 9.5, compounded semiannually
A = $

Answer by lwsshak3(6529)

Find the accumulated amount A if the principal P is invested at the interest rate of r/year for t yr. (Round your answer to the nearest cent.)
P = $1000, r = 4%, t = 9.5, compounded semiannually
**
compound interest formula: A=P(1+r)^t, P=initial investment, r=interest rate/period, A=amount after t periods
For given problem:
r=.04/2=.02 (interest rate/period)
t=19.5*2=39 periods
A=1000(1+.02)^39
A=1000(1.02)^39
A=1000*2.1647=$2164.70